1/3

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

The Werewolf The Bans- hee (An Ap- proach) Le loup-garou Varulven

tradukita de Alexander Gross tradukita de Max Knight tradukita de R. Platteau tradukita de Sven Collberg

A Werewolf, troubled by

his name, Left wife and brood one

night and came To a hidden graveyard

to enlist The aid of a long-dead

philologist.

One night, a banshee

slunk away from mate and child,

and in the gloom went to a village te-

acher’s tomb, requesting him: “In-

flect me, pray.”

Un loup-garou, une cer-

taine nuit, Laissa sa moitié et puis se

rendit Auprès de la tombe d’un

maître d’école Et lui dit, “ sur moi dis

tout, c’est ton rôle ! ”

En varulv lopp

från fru m.m. en vacker afton

samt begav sig till en bymagi-

sters grav och bad: “Var

snäll och konjugera.

“Oh sage, wake up, plea-

se don’t berate me,” He howled sadly, “Just

conjugate me.” The seer arose a bit un-

steady Yawned twice, wheezed

once, and then was ready.

The village teacher climbed up straight

upon his grave stone

with its plate and to the appariti-

on said who meekly knelt

before the dead:

Un loup-garou, une cer-

taine nuit, Laissa sa moitié et puis se

rendit Auprès de la tombe d’un

maître d’école Et lui dit, “ sur moi dis

tout, c’est ton rôle ! ” mig!” Bys-

kolläraren stod upp

på blecknamnsskyltens mässingknopp

med svar till den, som satt på pass

med tåligt korslagd tass på tass.

“Well, ‘Werewolf’ is

your plural past, While ‘Waswolf’ is sin-

gularly cast: There’s ‘Amwolf’ too,

the present tense, And ‘Iswolf,’ ‘Arewolf’

in this same sense.”

“The banSHEE, in the subject’s place;

the banHERS, the

possessive case. The banHER, next,

is what they call objective case–and

that is all.”

“ - On peut à ton sujet

puisque nous y sommes Se demander encor bien

des choses en somme. Loup-garoù ? mais avant Il faut se dire : loup-

garquand ?

“Jag var-ulv 1:a

pers. i sing., du, han, hon, den,

det var-ulv. Så vi voro-ulv, I vor-

en. På de voro-ulv kom

ingenting.”

…  …  …  …

2/3

“I know that–I’m no

mental cripple– The future form and

participle Are what I crave,” the

beast replied. The scholar paused–

again he tried:

The banshee marve-

led at the cases and writhed with

pleasure, making faces,

but said: “You did

not add, so far, the plural to the sin-

gular!”

Loup-garcomment ? et

puis Bien évidemment loup-

garqui ? ” Toutes ces questions sem-

blèrent au loup bien absconses,

Car en fait ce qu’il voulait c’était les réponses.

Varulven myste

gladåt hur pers. hade följts

av sing. och plur.

“Fast”, bad han, “kan det inte ges ens

en enda pers. i tempus presens?”

“A ‘Will-be-wolf?’ It’s

just too long: ‘Shall-be-wolf?’ ‘Has-

been-wolf?’ Utterly wrong!

Such words are wounds

beyond all suture– I’m sorry, but you have

no future.”

The teacher, though,

admitted then that this was not wi-

thin his ken. “While bans are fre-

quent”, he advised,

“A she cannot be plurized.”

“ - Point de réponses, dit le

puits de savoir, Etre imaginaire, nul ne

peut te voir ! ” A ces mots l’animal éclata

tout en larmes. “ - Oh!, même les baisers

de ma douce moitié Bekänna måste

vår magister, att därvidlag

fanns vissa brister.

En varulv sine im-

perfecto var contradictio in

adjecto.

The Werewolf knew better–his sons still slept

At home, and home-

wards now he crept, Happy, humble, without

apology For such folly of philolo-

gy.

The banshee, rising

clammily, wailed: “What about

my family?” Then, being not a

learned creature, said humbly

“Thanks” and left the teacher.

Ne pourrons plus me faire

me sentir tout entier. Rester un loup-garou est

un sort qui m’alarme. ” Mais comme le loup était

bien poli Il remercia et puis repartit.

Varulven satt med

tårad lins. Han visste att

han fanns och finns.

Därtill kom också

fru m.m. som även måste

existera.

…  …  …  …

3/3

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu  ĉi  angli-
go  troviĝas  en

http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_ url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Francan de R. Platteau.

Arg-825-1655 (2013-02-13 21:56:38)

Tiu ĉi francigo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Svedan de Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .